Oakdale Fashions, Inc.’s, 2021 income statement is reported below. Oakdale Fashions, Inc., Income Statement for 2021

Net sales (all credit) $ 565,000

Less: Cost of goods sold 215,000

Gross profits $ 350,000

Less: Other operating expenses 90,000

EBITDA $ 260,000

Less: Depreciation and amortization 15,000

EBIT $ 245,000

Less: Interest 80,000

EBT $ 165,000 Less:

Taxes Net income $ Determine the firm's 2021 tax liability. (Round your answer to the nearest dollar amount.) Determine the firm's 2021 net income. (Round your answer to the nearest dollar amount.)
Answer #1

Post December 31, 2017 the corporate tax rate is changed to 21% in US. So assuming that the company has its operations in USA, taxes payable is calculated as EBT * Tax rate

So Tax Payable = 165000 * 21% = $34,650

Net Income = EBT - Tax payable = 165000 - 34650= $130,350